OpenStudy (anonymous):

given the differential dy/dx=(x+y+1)/(x+y+3). by using the substitution z=x+y+1,find the solution of the equation

OpenStudy (anonymous):

Replace x+y by z-1 up and down Then continue :)

OpenStudy (unklerhaukus):

what is dz/dx?

OpenStudy (anonymous):

1+(dy/dx)

OpenStudy (anonymous):

btw this question under topic of ODE of first order

OpenStudy (unklerhaukus):

so dy/dx = dz/dx - 1

OpenStudy (unklerhaukus):

What is the equation now? (in terms on dz/dx, and z)

OpenStudy (unklerhaukus):

\[\frac{dy}{dx}=\frac{x+y+1}{x+y+3}\] letting \(z=x+y+1\); \[\dfrac{dz}{dx}=1+\dfrac{dy}{dx}\\\dfrac{dy}{dx}=\dfrac{dz}{dx}-1\] Substituting these into the DE \[\dfrac{dz}{dx}-1=\frac{z}{z+2}\]